OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?

OpenStudy (anonymous):

what kind of a question is this?

OpenStudy (anonymous):

is it 450, 520, 900, or 620?

sam (.sam.):

Let A be the number of one-step equations eliminated and B the number of two-step equations eliminated then, A + B = 1070 .....Equation (1) and we are told that, 3A - 2B = 110.....Equation (2) multiplying Eqn (1) by 2 gives 2A + 2B = 2140........now add Eqn(2) 3A - 2B = 110 5A = 2250 , A = 450

OpenStudy (ragingsquirrel):

Math Idol.... whoever came up with that must be pretty bored